OpenStudy (anonymous):

What is the value of the discriminant for the function y=3x^2+4x-5

OpenStudy (aravindg):

discriminant =b^2-4ac

OpenStudy (aravindg):

here d=16-4*3*-5 =76

OpenStudy (aravindg):

gt it>?

OpenStudy (anonymous):

wow thats it? 76 is the anwser

OpenStudy (aravindg):

yep